Please answer all parts
Given Data: AUBURN CAR WASH Cost of new car wash machines Estimated residual value Machine life in years Machine life in hours $ 500,000 $20,000 10,000 Actual use per year: Hours Used Year 2,000 1,500 1,400 1,800 2,000 1,300
media%2F348%2F34838f01-c065-4763-a936-78
media%2F4df%2F4df1bc64-83ab-4807-a509-3c
media%2F1b4%2F1b4e48de-2b9f-429f-bc81-10
Answer #1
Auburn Car Wash ( Activity Activity Based)
Calculation = End of Year Amount
Year Hour used
(i)
Depericaition rate(2) Deperication expense
3=(1*2)     
Accumulated depericiation Book value ( Cost of machine-Accum deperication  
1 2000 48 $   96,000.00 $          96,000.00 $           404,000.00
2 1500 48 $   72,000.00 $        168,000.00 $           332,000.00
3 1400 48 $   67,200.00 $        235,200.00 $           264,800.00
4 1800 48 $   86,400.00 $        321,600.00 $           178,400.00
5 2000 48 $   96,000.00 $        417,600.00 $             82,400.00
6 1300 48 $   62,400.00 $        480,000.00 $             20,000.00
Refer
Depericiation charges = (Cost of Machine- Salvage Value)
Total Estimated Hours
= (500000-20000)
10000
= 48
Auburn Car Wash ( Straight Line)
Calculation = End of Year Amount
Year Depriciable cost (1) Depericaition rate (2) Deperication expense
3=(1*2)     
Accumulated depericiation Book value ( Cost of machine-Accum deperication  
1 $          500,000 16.67% $   83,350.00 $          83,350.00 $           416,650.00
2 $          416,650 16.67% $   69,455.56 $        152,805.56 $           347,194.45
3 $          347,194 16.67% $   57,877.31 $        210,682.87 $           289,317.13
4 $          289,317 16.67% $   48,229.17 $        258,912.03 $           241,087.97
5 $          241,088 16.67% $   40,189.36 $        299,101.40 $           200,898.60
6 $          200,899 16.67% $   33,489.80 $        332,591.20 $           167,408.80
Refer
Depericiation rate = 1
Total estimated life
= 1 = 16.67%
6
Auburn Car Wash ( Double declining Balance)
Calculation = End of Year Amount
Year Beginning Boook value X Depericaition rate Deperication expense
  
Accumulated depericiation Book value ( Cost of machine-Accum deperication   Refer for depericiation Expense
1 $          500,000 0.33 $      165,000 $              165,000 $                 335,000 500000*0.33
2 $          335,000 0.33 $      110,550 $              275,550 $                 224,450 ( 500000-165000)*0.33)
3 $          224,450 0.33 $         74,069 $              349,619 $                 150,382 ( 500000-275500)*0.33)
4 $          150,382 0.33 $         49,626 $              399,244 $                 100,756 ( 500000-349619)*0.33)
5 $          100,756 0.33 $         33,249 $              432,494 $                   67,506 ( 500000-399244)*0.33)
6 $            67,506 0.33 $         22,277 $              454,771 $                   45,229 ( 500000-432494)*0.33)
Refer
Depericiation rate = 1 *2
Total estimated life
= 1 *2= 16.67%
6

It is observed that at the end of 6th year book value will come higher by calculating a depreciation through straight line value method